# Robust Denoising Neural Reranker for Recommender Systems

### v1

@@ 4.4 Sensitivity Analysis (RQ6)
We conduct sensitivity analysis for $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$ for both DNR-G and DNR-B, and present the results in Figure 3. While the Beta noise generally performs better than Gaussian noise in most settings, there exists some common patterns in these hyper-parameters: 1) all parameters appears to have an optimal value within the searched region, indicating the effectiveness of including each corresponding designs, i.e., $\lambda_{c}$ for the posterior design Eq.(4), $\lambda_{m}$ for the combined reconstruction loss Eq.(6), and $\lambda_{e}$ for the switching from heuristic sampler to model-based sampler; 2) over amplify these hyper-parameters results in the downgrade of the rerank performance, indicating the significance of keeping a balance between the involved components. Additionally, the analysis of $\lambda_{c}$ in the left plot of Figure 3 shows that using the model-based noise sampler can further enhance the performance over the heuristic noise sampler.
Figure 3: Sensitivity analysis of our method on Kuaivideo dataset.
@@ 5 Conclusion
In this paper, we provide evidence that the initial retriever scores are informative for the latter reranker stage in the multi-stage recommender system. To fully utilize such scores, we formulate the reranker as a noise reduction problem in the item scoring space and propose DNR, which collaboratively learns a more robust denoising reranker and a generative noise injection posterior. Theoretical analysis and empirical validation across three public datasets demonstrate the effectiveness of the noise reduction formulation and the design of our DNR. We also show that the posterior is intuitively and empirically a beta distribution, which is potentially more suitable than Gaussian noise in recommendation tasks. While a heuristic design of the noisy posterior can effectively improve the model performance, we prove that we can more comprehensively optimize a learnable posterior with the theoretically derived learning objective, and further boost the reranker’s performance.
Appendix A Proofs
A.1 Derivation of the Three Objectives in our Solution
In this section, we illustrate the derivation of the Eq.(5). Similar to the variational Bayesian inference [32], we start by analyzing the KL divergence between the ground truth and estimated posterior:
@@ A.1 Derivation of the Three Objectives in our Solution
Note that a key difference between Eq.(9) and the conventional derivation in [32] is that the variational posterior $\phi$ is not a latent factor encoder that reduces noise and dimensions, but an explicit noise injection model that mimics the behavior of retriever scores. In terms of the denoising problem formulation, it is also similar to a one-step diffusion, but with a key difference in that the ground truth $\mathbf{x}_{u}$ is not sampled by a manually defined distribution. Instead, an existing retriever $p_{\text{x}}$ serves as the prior.
@@
$=$ $-\mathbb{E}_{p_{\phi}}\big[\log\frac{p(\mathbf{x},\mathbf{z}_{u})}{p_{\phi}(\mathbf{x}|\mathbf{z}_{u})}\big]-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)$
$=$ $-\mathbb{E}_{p_{\phi}}\big[\log p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})\big]+D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x}}(\mathbf{x})\big)-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)$
$=$ $-\mathbb{E}_{p_{\phi}}\big[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})\big]+\mathbb{E}_{p_{\phi}}\big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})}\big]+D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x}}(\mathbf{x})\big)$
@@
which is equivalent to Eq.(5). It generally follows the derivation of the evidence lower bound, but we further decompose the reconstruction term into the parameterized reconstruction loss $\mathcal{L}_{\text{z}}$ and the adversarial posterior learning loss $\mathcal{L}_{\text{adv}}$. This decomposition indicates that the reranker $q_{\theta}$ is simultaneously an approximation of the ground truth $p_{\text{z|x}}$ as well as a premise factor that determines $p_{\text{z|x}}$, and their difference mainly describes the reranker’s error in predicting user feedback.
A.2 The Probabilistic View of Standard Reranking Models
In this section, we formally analyze the Eq.(3). Recall that the direct optimization goal $\mathcal{L}_{\text{std}}=-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})]$, where the observed $\mathbf{x}_{u}$ only comes from the retriever prior $p_{\text{x}}$ without posterior modeling. Then, we can extract the error of $q_{\theta}$ and derive the following relation:
@@ A.2 The Probabilistic View of Standard Reranking Models
$-D_{\text{KL}}\big(p_{x}(\mathbf{x}_{u})\|p_{\text{x|z}}(\mathbf{x}_{u}|\mathbf{z}_{u})\big).$
@@
As we have illustrated in section 3.2, the direct optimization of the first term $\mathcal{L}_{\text{direct}}$ might be misaligned with the data likelihood maximization goal when there exists irregular retriever scores or the posterior is inconsistent with the prior distribution. In contrast, our proposed solution solves this limitation by collaboratively learning the noise-injecting posterior $p_{\phi}$ along with the reranker. The resulting objectives in Eq.(5) intentionally finds a posterior that align with the ground truth prior and explore adversarial cases that the reranker fail to accurately reconstruct the user feedback (details in section 3.4).
Appendix B Learning Algorithm
@@
0: Hyperparameters $\lambda_{c}$ (or $\lambda_{e}$ if using learnable posterior), $\lambda_{m}$
0: Optimized denoising model $q_{\theta}$ (and the posterior $p_{\phi}$ if using learnable posterior).
@@
2:  if Using learnable posterior and $i>\lambda_{e}$ then
@@
4:    Sample $\mathbf{x}\sim p_{\phi}^{\text{model}}$ for given $\mathbf{u}$ and $\mathbf{z}_{u}$;
5:    Optimize $\theta$ based on $\mathcal{L}_{\theta}$;
6:    Optimize $\phi$ based on $\mathcal{L}_{\text{adv}}$ and $\mathcal{L}_{\text{x}}^{\prime}$;
@@
10:    Sample $\mathbf{x}\sim p_{\phi}^{\text{heuristic}}$ for given $\mathbf{u}$ and $\mathbf{z}_{u}$
11:   end for
12:  end if
13:end for
@@
Appendix C Detailed Settings for Offline Experiments
C.1 Datasets
The ML-1M dataset is a public benchmark in the field of recommender systems, comprising 1 million ratings from 6,040 users across more than 3,900 movies. The Kuaivideo dataset is sourced from Kuaishou, a popular short-video app with over 300 million active users daily. The Amazon Books dataset is a comprehensive collection of product reviews specifically focused on books available on Amazon. For all datasets, we initially process by eliminating users and items with fewer than 20 interactions, to avoid the cold-start issue. For the retriever stage, we split the data into train and test subsets at a ratio of 8:2. For the reranker stage, we sort the interactions in chronological order, using the last six interactions as the item list revealed to users after reranking. We present the statistics of the processed datasets for the retriever and reranker stages in Table LABEL:dataset_stage1 and Table LABEL:dataset_stage2, respectively.
@@ C.1 Datasets
C.2 Baselines
@@
  • SASRec [23] proposes a self-attention based sequential recommendation model (SASRec) that captures long-term semantics and short-term dynamics using an attention mechanism.
  • Caser [36] introduces a convolutional sequence embedding recommendation model that captures sequential patterns using convolutional filters applied to latent space embeddings.
  • GRU4Rec [37] apply recurrent neural networks (RNNs) to session-based recommendations, addressing the challenge of short user sessions without long-term profiles.
  • MiDNN [24] proposes a global optimization framework for e-commerce search ranking that considers mutual influences between items, using extended features and sequence generation to optimize ranking and maximize overall utility.
@@
  • DLCM [8] leverages local ranking context from top-retrieved documents to refine initial ranking lists. The model uses a recurrent neural network to capture document interactions and an attention-based loss function, significantly improving performance over traditional learning-to-rank methods.
  • SetRank [38] introduces a neural learning-to-rank model that learns a permutation-invariant ranking function directly on document sets. Utilizing multi-head self-attention blocks, SetRank captures cross-document interactions and achieves robust performance across varying input sizes.
  • PRM [6] addresses personalized re-ranking in recommendation systems by incorporating user-specific preferences into the re-ranking process. It considers user behavior and candidate list, enhancing both relevance and personalization in the final recommendations.
  • MIR [25] proposes to capture complex interactions between user actions and candidate list features across multiple levels, improving the accuracy and relevance of recommendations through a hierarchical reranking approach.
@@
  • EGRerank [22] proposes an evaluator-generator framework for learning-to-rank (LTR) in e-commerce applications. The framework includes an evaluator that evaluates recommendations involving context, a generator that maximizes the evaluator score through reinforcement learning, and a discriminator that ensures the generalization of the evaluator.
  • Pier [9] follows a two-stage architecture with a Fine-grained Permutation Selection Module (FPSM) and an Omnidirectional Context-aware Prediction Module (OCPM). FPSM selects top-K candidate permutations based on user interest using SimHash, while OCPM evaluates these permutations with an omnidirectional attention mechanism.
  • NAR4Rec [10] explores the use of non-autoregressive generative models for re-ranking in recommendation systems, addressing challenges such as sparse training samples and dynamic candidates by introducing a matching model, unlikelihood training to distinguish feasible sequences, and contrastive decoding to capture item dependencies.
@@
  • DiffuRec [26] incorporates a diffusion process to inject noise into target item embeddings and a reverse process to reconstruct the target item representation.
  • DCDR [11] introduces a new framework that leverages diffusion models for the reranking stage in recommendation systems. DCDR extends traditional diffusion models with a discrete forward process and a conditional reverse process to generate high-quality item sequences.
C.3 Implementation of Diverse Strategies for Leveraging Retriever Scores
Here, we present the detailed implementation of different ways to leverage retriever scores in Figure C.3. The left one refers to “c score”, which leverages a concatenate operation to combine the score features with encoded user states from users’ interaction history. The moderate one represents “+ score”, which incorporates the score feature with user states by a simple plus operation. The right one is the “w score”, which adopts the initial candidate score as the weight of candidate items.s

### v2

@@ 4.4 Sensitivity Analysis (RQ6)
We conduct sensitivity analysis for $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$ for both DNR-G and DNR-B, and present the results in Figure 3. While the Beta noise generally performs better than Gaussian noise in most settings, as shown in Table 2, there exist some common patterns in these hyper-parameters: 1) all parameters appears to have an optimal value within the searched region, indicating the effectiveness of including each corresponding designs, i.e., $\lambda_{c}$ for the synthetic noisy scores in Eq.(4), $\lambda_{m}$ for the combined denoising loss Eq.(6), and $\lambda_{e}$ for the switching from heuristic noise generator to model-based one; 2) Over-amplifying these hyper-parameters results in the downgrade of the rerank performance, indicating the significance of keeping a balance between the involved components. Additionally, the analysis of $\lambda_{c}$ in the left plot of Figure 3 shows that using the model-based noise generator can further enhance the performance over the heuristic one.
Figure 3: Sensitivity analysis of hyperparameters (i.e., $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$) for our method on the Kuaivideo dataset.
@@ 5 Conclusion
In this paper, we provide evidence that the initial retriever scores are informative for the latter reranker stage in the multi-stage recommender system. To formulate the relationship between retriever and reranker scores, we view the reranker as a noise reduction task from the retriever scores and propose DNR. To fully explore the score space, we propose the framework of DNR, adding noise to generate retriever scores as augmentation and then denoising them to align with user feedback. While a heuristic design of the noise distribution can effectively improve the model performance (e.g., Gaussian or Beta distribution), we prove that adversarially learns a denoising reranker and a noise generator for the retriever stage can further boost the reranker’s performance. Theoretical analysis and empirical validation across three public datasets demonstrate the effectiveness of the noise reduction formulation and the design of our DNR.
Ethics Statement.
This work is designed to explore the significance of retriever scores for the reranker, formulating the reranker as a denoising process of retriever scores with an adversarial framework. We do not foresee any direct, immediate, or negative societal impacts of our research.
Reproducibility Statement.
All the results in this work are reproducible. We have discussed the optimal hyperparameters and the details on devices and software environments in Section 4.1 and Appendix D.2.
Appendix A Notations
symbol description
$\mathbf{u}$ user request information including profile features and interaction history
$\mathbf{x}$ random variables of retriever scores on the candidate item set
$\mathbf{z}$ random variables of user feedback on the candidate item set
$\mathbf{x}_{u}$, $\mathbf{z}_{u}$ observed retriever scores and user feedback in data
$\mathbf{x}_{u}^{\prime}$ the generated retriever scores from the learned or heuristic posterior model $\phi$
$q_{\theta}$ the probabilistic representation of the reranker model
$p_{\phi}$ the probabilistic representation of the posterior
$f_{\phi}$ the noise generator that implements the posterior
$p_{\text{x}}$ the probabilistic representation of the non-optimizable retriever
$p_{\text{x|z}}$ the posterior of a given reranker
$p_{\text{z|x}}$ the probabilistic representation of ground truth user feedback
$\mathcal{L}_{\text{direct}}$ the standard user feedback alignment loss given the retriever scores
$\mathcal{L}_{\text{z}}$ the user feedback alignment loss given generated retriever scores from posterior
$\mathcal{L}_{\text{adv}}$ the adversarial loss that encourages the posterior to generate irregular noises
$\mathcal{L}_{\text{x}}$ the regularization of posterior
$\lambda_{c}$ the hyperparameter that controls the magnitude of noise injection in Eq.(4)
$\lambda_{m}$ the hyperparameter that balances $\mathcal{L}_{\text{direct}}$ and $\mathcal{L}_{z}$ when learning the reranker
$\lambda_{e}$ the hyperparameter that determines which epoch switches the noise generator
Table 4: List of Notations
Appendix B Proofs
B.1 Derivation of the Three Objectives in our Solution
In this section, we illustrate the derivation of the Eq.(5). Similar to the variational Bayesian inference Kingma & Welling (2014), we start by analyzing the KL divergence between the real retriever scores and synthetic noisy scores:
@@ B.1 Derivation of the Three Objectives in our Solution
Note that a key difference between Eq.(8) and the conventional derivation in Kingma & Welling (2014) is that the variational posterior $\phi$ is not a latent factor encoder that reduces noise and dimensions, but an explicit noise generator that mimics the behavior of retriever scores. In terms of the denoising problem formulation, it is also similar to a one-step diffusion, but with a key difference in that the real retriever scores $\mathbf{x}_{u}$ are not sampled by a manually defined distribution. Instead, an existing retriever $p_{\text{x}}$ serves as the prior distribution.
@@
$=$ $-\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\big[\log\frac{p(\mathbf{x},\mathbf{z}_{u})}{p_{\phi}(\mathbf{x}|\mathbf{z}_{u})}\big]-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)$
$=$ $-\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\big[\log p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})\big]+D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x}}(\mathbf{x})\big)-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)$
$=$ $-\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\big[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})\big]+\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})}\big]+D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x}}(\mathbf{x})\big)$
@@
which is equivalent to Eq.(5). It generally follows the derivation of the evidence lower bound in variational auto-encoders, except that we further decompose the first reconstruction term into the augmented denoising loss $\mathcal{L}_{\text{z}}=-\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\big[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})\big]$ and the adversarial learning loss $\mathcal{L}_{\text{adv}}=\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})}\big]$. Intuitively, when $q_{\theta}$ is closely aligned with the real user feedback distribution $p_{\text{z|x}}$, the adversarial term $\mathcal{L}_{\text{adv}}$ will be close to zero. Minimizing $\mathcal{L}_{\text{adv}}$ would encourage the noise generator $f_{\phi}$ to find samples where $q_{\theta}$ and $p_{\text{z|x}}$ behave differently.
B.2 The Analytical Limitation of Standard Reranking Models
In this section, we formally analyze the Eq.(3). Recall that the direct optimization goal $\mathcal{L}_{\text{direct}}|_{\mathbf{x}=\mathbf{x}_{u}}=-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u})]$, where the observed $\mathbf{x}_{u}$ only comes from the retriever prior $p_{\text{x}}$ without posterior modeling. Note that we additionally define $p_{\text{z|x}}$ as the ground-truth user feedback probability and $p_{\text{x|z}}$ as the posterior distribution of retriever scores conditioned on user feedback. Both $p_{\text{z|x}}$ and $p_{\text{x|z}}$ are not related to the reranker $q_{\theta}$, and they are merely determined by the retriever prior and the user. Then, we can extract the user behavior alignment error of $q_{\theta}$ as follows and derive its relation to the goal of $\max_{\theta}\log p(\mathbf{z}_{u})$:
@@ B.2 The Analytical Limitation of Standard Reranking Models
$-D_{\text{KL}}\big(p_{\text{x}}(\mathbf{x}_{u})\|p_{\text{x|z}}(\mathbf{x}_{u}|\mathbf{z}_{u})\big).$
@@
As we have discussed in section 3.2, the direct optimization of the first term $\mathcal{L}_{\text{direct}}$ might cause misalignment with the data likelihood maximization goal, since the observed data does not provide sufficient information about the retriever score distribution and the reranker is unaware of the noise patterns in the retriever score space. In contrast, our proposed solution solves this limitation by collaboratively learning the noise generator $f_{\phi}$ along with the reranker. The resulting objectives in Eq.(5) intentionally find a synthetic noisy score distribution $p_{\phi}$ that aligns with the real retriever score distribution and explores adversarial cases in which the reranker fails to accurately denoise into user feedback (details in Section 3.4).
Appendix C Learning Algorithm
@@
0: Hyperparameters $\lambda_{c}$, $\lambda_{e}$, and $\lambda_{m}$, initial reranker $\theta$ and noise generator $\phi$
0: Optimized denoising reranker $q_{\theta}$ and the noise generator $f_{\phi}$.
@@
2:  if $i>\lambda_{e}$ then
@@
4:    Generate $\bm{\epsilon}\sim f_{\phi}^{\text{model}}$ for given $\mathbf{u}$ and $\mathbf{z}_{u}$ and form synthetic noisy scores $\mathbf{x}_{u}^{\prime}$ with Eq (4)
5:    Optimize reranker $\theta$ based on $\mathcal{L}_{\theta}$;
6:    Optimize noise generator $\phi$ based on $\mathcal{L}_{\text{adv}}$ and $\mathcal{L}_{\text{x}}$;
@@
10:    Sample $\bm{\epsilon}\sim f_{\phi}^{\text{heuristic}}$ for given $\mathbf{u}$ and $\mathbf{z}_{u}$ and form synthetic noisy scores $\mathbf{x}_{u}^{\prime}$ with Eq (4)
11:    Optimize reranker $\theta$ based on $\mathcal{L}_{\theta}$;
12:   end for
13:  end if
14:end for
@@
Appendix D Detailed Settings for Offline Experiments
D.1 Datasets
The ML-1M dataset is a public benchmark in the field of recommender systems, comprising 1 million ratings from 6,040 users across more than 3,900 movies. The Kuaivideo dataset is sourced from Kuaishou, a popular short-video app with over 300 million active users daily. The Amazon Books dataset is a comprehensive collection of product reviews specifically focused on books available on Amazon. For all datasets, we initially process by eliminating users and items with fewer than 20 interactions, to avoid the cold-start issue. For the retriever stage, we split the data into train and test subsets at a ratio of 8:2, where each sample is composed of interaction history and an item for prediction. For the reranker stage, we sort the interactions in chronological order, using the last six interactions as the item list revealed to users after reranking. We present the statistics of the processed datasets for the retriever and reranker stages in Table LABEL:dataset_stage1 and Table LABEL:dataset_stage2, respectively.
@@ D.1 Datasets
D.2 Hyperparameters Setting
Here we present the optimal hyperparameters’ settings in Table 7.
Datasets ML-1M Kuaivideo Book
$\lambda_{c}$ 0.4 0.3 0.5
$\lambda_{m}$ 0.4 0.6 0.4
$\lambda_{e}$ 40 80 100
weight_decay 0.001 0.001 0
dropout 0.3 0.3 0.5
batch_size 2048 2048 2048
emb_dim 128 128 128
D.3 Baselines
@@
  • SASRec Kang & McAuley (2018) proposes a self-attention based sequential recommendation model (SASRec) that captures long-term semantics and short-term dynamics using an attention mechanism.
  • Caser Tang & Wang (2018) introduces a convolutional sequence embedding recommendation model that captures sequential patterns using convolutional filters applied to latent space embeddings.
  • GRU4Rec Hidasi et al. (2016) apply recurrent neural networks (RNNs) to session-based recommendations, addressing the challenge of short user sessions without long-term profiles.
  • MiDNN Zhuang et al. (2018) proposes a global optimization framework for e-commerce search ranking that considers mutual influences between items, using extended features and sequence generation to optimize ranking and maximize overall utility.
@@
  • DLCM Ai et al. (2018) leverages local ranking context from top-retrieved documents to refine initial ranking lists. The model uses a recurrent neural network to capture document interactions and an attention-based loss function, significantly improving performance over traditional learning-to-rank methods.
  • SetRank Pang et al. (2020) introduces a neural learning-to-rank model that learns a permutation-invariant ranking function directly on document sets. Utilizing multi-head self-attention blocks, SetRank captures cross-document interactions and achieves robust performance across varying input sizes.
  • PRM Pei et al. (2019) addresses personalized re-ranking in recommendation systems by incorporating user-specific preferences into the re-ranking process. It considers user behavior and candidate list, enhancing both relevance and personalization in the final recommendations.
  • MIR Xi et al. (2022) proposes to capture complex interactions between user actions and candidate list features across multiple levels, improving the accuracy and relevance of recommendations through a hierarchical reranking approach.
@@
  • EGRerank Huzhang et al. (2023) proposes an evaluator-generator framework for learning-to-rank (LTR) in e-commerce applications. The framework includes an evaluator that evaluates recommendations involving context, a generator that maximizes the evaluator score through reinforcement learning, and a discriminator that ensures the generalization of the evaluator.
  • Pier Shi et al. (2023) follows a two-stage architecture with a Fine-grained Permutation Selection Module (FPSM) and an Omnidirectional Context-aware Prediction Module (OCPM). FPSM selects top-K candidate permutations based on user interest using SimHash, while OCPM evaluates these permutations with an omnidirectional attention mechanism.
  • NAR4Rec Ren et al. (2024) explores the use of non-autoregressive generative models for re-ranking in recommendation systems, addressing challenges such as sparse training samples and dynamic candidates by introducing a matching model, unlikelihood training to distinguish feasible sequences, and contrastive decoding to capture item dependencies.
@@
  • DiffuRec Li et al. (2024) incorporates a diffusion process to inject noise into target item embeddings and a reverse process to reconstruct the target item representation.
  • DCDR Lin et al. (2024) introduces a new framework that leverages diffusion models for the reranking stage in recommendation systems. DCDR extends traditional diffusion models with a discrete forward process and a conditional reverse process to generate high-quality item sequences.
D.4 Implementation of Diverse Strategies for Leveraging Retriever Scores
Here, we present the detailed implementation of different ways to leverage retriever scores in Figure D.4. The left one refers to “c score”, which leverages a concatenate operation to combine the score features with encoded user states from users’ interaction history. The moderate one represents “+ score”, which incorporates the score feature with user states by a simple plus operation. The right one is the “w score”, which adopts the initial candidate score as the weight of candidate items.